Presented below are condensed financial statements adapted from those of two actual companies competing in the pharmaceutical industry—Johnson and Johnson (J&J) and Pfizer, Inc. ($ in millions, except per share amounts). Balance Sheets ($ in millions, except per share data) J&J Pfizer Assets: Cash $ 18,143 $ 12,398 Short-term investments 6,390 12,646 Accounts receivable (net) 9,184 11,385 Inventories 6,016 10,135 Other current assets 6,100 5,965 Current assets 45,833 52,529 Property, plant, and equipment (net) 16,254 24,695 Intangibles and other assets 20,030 73,355 Total assets $ 82,117 $ 150,579 Liabilities and Shareholders' Equity: Accounts payable $ 7,396 $ 5,031 Short-term notes 5,297 12,976 Other current liabilities 9,593 14,488 Current liabilities 22,286 32,495 Long-term debt 5,205 8,005 Other long-term liabilities 7,205 24,200 Total liabilities 34,696 64,700 Capital stock (par and additional paid-in capital) 5,820 69,750 Retained earnings 50,893 49,272 Accumulated other comprehensive income (loss) (950 ) 285 Less: Treasury stock and other equity adjustments (8,342 ) (33,428 ) Total shareholders' equity 47,421 85,879 Total liabilities and shareholders' equity $ 82,117 $ 150,579 Income Statements Net sales $ 51,492 $ 54,818 Cost of goods sold 14,444 12,100 Gross profit 37,048 42,718 Operating expenses 21,959 30,682 Other (income) expense—net (655 ) 3,880 Income before taxes 15,744 8,156 Tax expense 4,723 2,447 Net income $ 11,021 $ 5,709 * Basic net income per share $ 3.32 $ 0.48 * This is before income from discontinued operations. Evaluate and compare the two companies by responding to the following questions. Note: Because two-year comparative statements are not provided, you should use year-end balances in place of average balances as appropriate. Required: 1. For both companies, compute the ratios below. 2. Evaluate and compare the two companies.
Answer #1

1. Receivables Turnover = Net credit Sales / Average Accounts receivables

J&J = 44002 / 7154 = 6.15 times Pfizer = 47328 / 9355 = 5.06 times

2. Average Collection Period = 365 / Receivable turnover ratio

J&J = 365 / 6.15 = 59.35 days Pfizer = 365 / 5.06 = 72.13 days

3. Inventory Turnover = Cost of goods sold / Average Inventory

J&J = 12680 / 4112 = 3.08 times Pfizer = 10336 / 6751 = 1.53 times

4. Average Days in Inventory = 365 / Inventory turnover ratio

J&J = 365 / 3.08 = 118.51 days Pfizer = 365 / 1.53 = 238.56 days

5. Profit Margin = Net Income / Net Sales

J&J = 8061 / 44002 = 18.32% Pfizer = 3044 / 47328 = 6.43%

6. Asset Turnover = Net Sales / Total Assets

J&J = 44002 / 55775 = 78.89 times Pfizer = 47328 / 124287 = 38.08 times

7. Return on Assets = Net Income / Total Assets

J&J = 8061 / 55775 = 14.45% Pfizer = 3044 / 124287 = 2.45%

8. Equity Multiplier = Total Assets / Shareholders Equity

J&J = 55775 / 31425 = 1.77 Pfizer = 124287 / 69933 = 1.78

9. Return on Shareholders' Equity = Net Income / Shareholders Equity

J&J = 8061 / 31425 = 25.65% Pfizer = 3044 / 69933 = 4.35%

Which of the two companies appears more efficient in collecting its accounts receivable and managing its inventory? - J&J is efficient in collecting account receivable as its collection period is lesser than Pfizer (59 vs 72 days). J&J is efficient in manageing inventory as well (118 days vs 238 days)

Which of the two firms had greater earnings relative to resources available? - J&J has greater earnings with a ROA of 14.45%

From the perspective of a common shareholder, which of the two firms provided a greater rate of return? - J&J has higher return on Shareholders equity. We can also refer to basic net income per share in the Income statement.

From the perspective of a common shareholder, which of the two firms appears to be using leverage more effectively to provide a return to shareholders above the rate of return on assets? - Leverage is almost same for both i.e. equity multiplier but J&J is earning higher return than Pfizer, so it is using leverage more effectively.

Presented below are condensed financial statements adapted from those of two actual companies competing in the pharmaceutical35,023 Retained earnings Accumulated other comprehensive income (loss) Less: Treasury stock and other equity adjustments Tota* This is before income from discontinued operations. Evaluate and compare the two companies by responding to the following qAnalysis Which of the two companies appears more efficient in collecting its accounts receivable and managing its inventory?